OpenStudy (anonymous):

The area of a rectangle is given by A = x2 – 2x + 30. Find the value of x if the total area of the rectangle is equal to 45 square units.

OpenStudy (anonymous):

It would be x=5

jhonyy9 (jhonyy9):

so if you know that A=45 so than just make this equation equal with 45 and check it how what will get OK ?

jhonyy9 (jhonyy9):

so x^2 -2x +30=45

jhonyy9 (jhonyy9):

subtract from both sides 45 and will get x^2 -2x +30-45=45-45 x^2 -2x -15=0 (x-5)(x+3)=0 x-5=0 x=5 so just this value will be right because for x+3=0 you will get x=-3 but you know sure that a length of a side not can be negative never ... good luck hope so much that is understandably sure right bye